Tim (type B-positive blood) and Tina (A-negative blood) decide to have children. Both of their dads are AB-negative and their mothers O-positive.

a) What is the chance that they will have a child that is O-negative, a universal donor?

1/16

1/8

1/4

1/2

3/4

b) What is the chance that if they have two children that they will have the same blood type?

1/64

1/32

1/16

1/8

1/2

3/4
Answer #1

Answer:

a). 1/8

Explanation:

Based on the information, the followings are the genotypes of parents.

Tim (IBIO; Rh+Rh-) x Tina (IAIO; Rh-Rh-)---- parents

IAIO x IBIO = IAIO (A-1/4), IBIO (B-1/4), IAIB (AB-1/4)& IOIO(O-1/4)

Rh+Rh- x Rh-Rh- = Rh+Rh-(1/2)&Rh-Rh-(1/2)

Chance of O negative = IOIO; Rh-Rh- = 1/4*1/2= 1/8

B). 1/64

Explanation:

The frequency of any blood type =1/8

Therefore, ..

the chance that if they have two children that they will have the same blood type = 1/8*1/8=1/64